40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 14 23:14:12 2019, gautam, Update, CDS, c1rfm errors 
    Reply  Fri Feb 15 15:22:08 2019, gautam, Update, CDS, c1rfm errors persist Screenshot_from_2019-02-15_15-21-47.png
Message ID: 14457     Entry time: Fri Feb 15 15:22:08 2019     In reply to: 14455
Author: gautam 
Type: Update 
Category: CDS 
Subject: c1rfm errors persist 

I restarted c1scyc1rfm (so both sender and receiver models were cycled) and power-cycled the c1iscey and c1sus machines. The TRY PD is certainly seeing light - it is just not getting piped over to c1rfm. dmesg doesn't give any clues. I'm out of ideas.

P.S. The new reality seems to be that getting ITMY stuck in the event of a c1susaux reboot is inevitable. As is the practise for ITMX, I tried slowly ramping the PIT and YAW biases to 0 slowly - but in the process of ramping YAW to 0, the optic got stuck. I am ramping in steps of 0.1 (in units of the PIT/YAW sliders, waiting ~3 seconds between steps), I guess I can try ramping even more slowly.

Update: I power cycled the physical RFM switch. This necessitated reboot of all vertex FEs. But seems like things are back to normal now...

Note: to unstick ITMY, seems like the best approach is:

  1. Jiggle bias until SIDE shadow sensor is on average above it's half-light level. This is the critical step. A bias of +20000 cts on the fast SIDE output seems to help.
  2. Set YAW bias to -10, ramp down the BIAS in steps of 0.1, watching shadow sensor levels to ensure optic doesn't get stuck again.
  3. Hope for the best. Iterate if necessary.
Quote:

The pressure is still 2e-4 torr according to CC1 so I thought I'd give ASS debugging a go tonight. But the arm transmission signal isn't coming through to the LSC model from the end PDs - so a resurfacing of this problem. Rebooting the sender model, c1scy, did not fix the problem. Moreover, c1susaux is dead. The last time I rebooted it, ITMY got stuck so I'm not going to attempt a revival tonight.

Attachment 1: Screenshot_from_2019-02-15_15-21-47.png  30 kB  | Hide | Hide all
Screenshot_from_2019-02-15_15-21-47.png
ELOG V3.1.3-